OpenStudy (anonymous):

A conveyor belt carries supplies from the first floor to the second floor, which is 27 feet higher. The belt makes a 60° angle with the ground. How far do the supplies travel from one end of the conveyor belt to the other? Round your answer to the nearest foot. If the belt moves at 75 ft/min, how long, to the nearest tenth of a minute, does it take the supplies to move to the second floor?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

In a 30-60-90 triangle, the ratio of the lengths of the sides is: 1 : sqrt(3) : 2

OpenStudy (anonymous):

just do sin(60)=27/x

OpenStudy (anonymous):

multiply both sides by x then divide both sides by sin(60) to get your answer

OpenStudy (anonymous):

then take that answer and divide it by 75

OpenStudy (mathstudent55):

Here, the 27 ft side corresponds to the sqrt(3) side, and the unknown side corresponds to 2 27/sqrt(3) = x/2

OpenStudy (anonymous):

how would i put that into the calculator??

OpenStudy (mathstudent55):

You solve the proportion by cross multiplying: 27/sqrt(3) = x/2 xsqrt(3) = 2*27 x = 2*27/sqrt(3) 2 times 27 divided by sqrt of 3

OpenStudy (anonymous):

31?

OpenStudy (anonymous):

is it 31?

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

thank you :)